if it cost $78.86 to buy 3 tickets in japan plus 2 tickets in switzerlan. 3 tickets in switzerland plus 2 tickets in japan cost $74.84. how much does an average movie ticket cost in each country?
Answer by sophxmai(62)

x=Japan
y=Switzerland

3x+2y=78.86
2x+3y=74.84
Find x and y

Multiply each equation to cancel one of the variables out.
3x+2y=78.86 x2
6x+4y=157.72

2x+3y=74.84 x3
6x+9y=224.52

Now the two equations have 6x, and if you subtract one from the other, you can cancel out one of the variables.
6x+4y=157.72
6x+9y=224.52

-5y=-66.8
y=-66.8/-5
y=13.36

Now you have y, substitute y into an equation to find x.
2x+3y=74.84
2x+3(13.36)=74.84
2x+40.08=74.84
2x=34.76
x=34.76/2
x=17.38

So tickets in Japan cost $17.38 and tickets in Switzerland cost $13.36